OpenStudy (driftracer305):

solve the quadratic equation by completing the square. -3x^2+9x=1

OpenStudy (zzr0ck3r):

for the equation \(ax^2+bx+c=0\) we can complete the square with this \(a(x+\frac{b}{2a})^2=-c+a(\frac{b}{2a})^2\)

OpenStudy (driftracer305):

right..........but i did it in the classic way......jus got stuck in the options though...........so far i solved up till here......... -3x^2 +9x +81/4 - 1 - 81/4=0

OpenStudy (driftracer305):

confused bro..? i did the (b/2)^2

OpenStudy (driftracer305):

that gave me 81/4

OpenStudy (anonymous):

\[ -3x^2+9x = -3(\color{red}{x^2}\color{blue}{-3}\color{red}{x}) \]Now we use \[ (x+a)^2 = x^2+2a+a^2 \\ (\color{purple}{x+a})^2-\color{green}{a^2}=\color{red}{x^2}\color{blue}{+2a}\color{red}{x} \]We let \(-3=2a\) and get \(a=-3/2\) and \(a^2 = 9/4\).\[ x^2-3x = (\color{purple}{x-3/2})^2-\color{green}{9/4} \]

OpenStudy (anonymous):

We sub that back in \[ -3x^2+9x = -3[(x-3/2)^2-9/4] = -3(x-3/2)^2 - 27/4 \]

OpenStudy (anonymous):

Actually, probably should have stuck with\[ (x-3/2)^2-9/4 = -\frac 13 \]

OpenStudy (anonymous):

I think you can do the rest.

OpenStudy (driftracer305):

yea.....i see........but take a look at this please

OpenStudy (anonymous):

And?

OpenStudy (anonymous):

You have to finish the rest of it.

OpenStudy (anonymous):

You're not sure what to do next?

OpenStudy (driftracer305):

thats the problem dude.......i end up using calculator and dividing everything resulting in decimals

OpenStudy (anonymous):

Do you know how to add fractions? Don't use a calculator.

OpenStudy (driftracer305):

yes i do

OpenStudy (driftracer305):

i tried ............2 full pages sitting in trash.....so i came here

OpenStudy (driftracer305):

please help me man..............my brain is frozen......blank!?!?!

OpenStudy (anonymous):

So you want me to teach you how to add fractions?

OpenStudy (anonymous):

\[ (x-3/2)^2 = -\frac 13 + \frac 94 \]

OpenStudy (driftracer305):

lol......no wait

OpenStudy (driftracer305):

x - 3/2 = sq.root of 23/12........right?

OpenStudy (driftracer305):

k wat next now.....?

OpenStudy (driftracer305):

uhh @wio ??

OpenStudy (anonymous):

well it is \[ |x-3/2| =\sqrt{\frac{23}{12}} \]

OpenStudy (anonymous):

\[ x=\frac 32 \pm \sqrt{\frac{23}{12}} \]

OpenStudy (driftracer305):

yea.....exactly wat i said.....wat next though?

OpenStudy (driftracer305):

the answer choices dont quite match...........the sq.root portion.....

OpenStudy (driftracer305):

@wio ........u there buddy?

OpenStudy (driftracer305):

@zzr0ck3r .......u there?

OpenStudy (anonymous):

hold on

OpenStudy (driftracer305):

k sorry.........m hasty.....lol

OpenStudy (zzr0ck3r):

lol

OpenStudy (driftracer305):

@zzr0ck3r .........we are stuck here....nd u think it's funny?

OpenStudy (zzr0ck3r):

wio aint stuck, im guessing hes busy

OpenStudy (driftracer305):

dont think so......he got the first part right........but the 2nd one....he might be working on i guess?

OpenStudy (zzr0ck3r):

\( -3x^2+9x=1\implies 3x^2-9x=-1\implies3x^2-9x+1=0 \) so \(3(x^2-3x)+1=0\\3(x-\frac{3}{2})^2+1=3(\frac{3}{2})^2\\3(x-\frac{3}{2})^2=\frac{27}{4}-1=\frac{23}{4}\) so \((x-\frac{3}{2})^2=\frac{23}{12}\implies x-\frac{3}{2}=\pm\sqrt{\frac{23}{12}}\\x=\frac{3}{2}\pm \sqrt{\frac{23}{12}}\)

OpenStudy (driftracer305):

@zzr0ck3r ..... we got there buddy.........but take a look at this attachment please....

OpenStudy (zzr0ck3r):

D

OpenStudy (driftracer305):

@zzr0ck3r ........so what do you think..??.............

OpenStudy (zzr0ck3r):

\(x=\frac{3}{2}\pm \sqrt{\frac{23}{12}}\rightarrow\frac{3}{2}\pm\frac{\sqrt{3}}{\sqrt{3}}\sqrt{\frac{23}{12}}=\frac{3}{2}\pm\sqrt{\frac{69}{36}}=\frac{3}{2}\pm\frac{\sqrt{69}}{6}\)

OpenStudy (zzr0ck3r):

so D is the answer

OpenStudy (anonymous):

\[\sqrt{\frac{ 23 }{ 12 }(\frac{ 3 }{ 3 })} = \sqrt{\frac{ 69 }{ 36 }} = \frac{ \sqrt{69} }{ 6 }\]

OpenStudy (zzr0ck3r):

\(x=\frac{3}{2}\pm \sqrt{\frac{23}{12}}\rightarrow\frac{3}{2}\pm1*\sqrt{\frac{23}{12}}=\frac{3}{2}\pm\frac{\sqrt{3}}{\sqrt{3}}\sqrt{\frac{23}{12}}=\frac{3}{2}\pm\sqrt{\frac{69}{36}}=\frac{3}{2}\pm\frac{\sqrt{69}}{6}\)

OpenStudy (driftracer305):

ohh wow.........@zzr0ck3r and @Johnbc .....thanks guyz....nd @wio too..!!

OpenStudy (driftracer305):

so how did you get 3/3.....?

OpenStudy (zzr0ck3r):

look at my steps

OpenStudy (zzr0ck3r):

1 = sqrt(3)/sqrt(3)

OpenStudy (zzr0ck3r):

when we move that in the sqrt we get 3/3

OpenStudy (anonymous):

Its just multiplying the square root fraction by one as zzr0ck3r demonstrated

OpenStudy (driftracer305):

ahh...........i see......i didnt know u could substitute something in a sq.root like this....

OpenStudy (zzr0ck3r):

\(\frac{3}{2}\pm\color{blue}{1}*\sqrt{\frac{23}{12}}=\frac{3}{2}\pm\color{blue}{\frac{\sqrt{3}}{\sqrt{3}}}\sqrt{\frac{23}{12}}=\frac{3}{2}\pm\sqrt{\frac{23*\color{blue}{3}}{12*\color{blue}{3}}}=\frac{3}{2}\pm\frac{\sqrt{69}}{6}\)

OpenStudy (zzr0ck3r):

yeah \(\sqrt{a}*\sqrt{b} = \sqrt{ab}\)

OpenStudy (zzr0ck3r):

but \(\sqrt{a}+\sqrt{b}\ne \sqrt{a+b}\)

OpenStudy (zzr0ck3r):

in general....

OpenStudy (driftracer305):

k .... @zzr0ck3r ........ thx a lot man........ nd for the extra hints ....... appreciate it :D

OpenStudy (zzr0ck3r):

np, I guarantee you that wio was just busy:)

OpenStudy (driftracer305):

lol........ seems like u knw him well